OpenStudy (anonymous):

(2/m-4)-((4m-8)/m^2-2m-8)=(2/m+2) Solve the rational equation for m.

OpenStudy (anonymous):

\[\frac{ 2 }{ m-4 }-\frac{ (4m-8) }{ m^2-2m-8 }=\frac{ 2 }{ m+2 }\]